# Universal Analog Computation: Fraïssé limits of dynamical systems

### v1

@@ 7. Proof of the algebroidality theorem
The remaining part (2) of being semi-algebroidal we show in two steps. First, by showing it when understanding ‘finite’ in the set-theoretic sense, and then showing that this is equivalent to the category-theoretic sense.
We start, though, with a lemma, which is a version of [57, lem. 2.1]. For completeness, we add a proof in appendix A.4.
@@
To show the equivalence of the two senses of ‘finite’, we need one more standard lemma on the interplay between partitions and projections in inverse limits. For completeness, we add a proof in appendix A.4.
@@ 8. The universal and homogeneous nondeterministic system
Now we can show the universality result for $\mathsf{Sys}^{\mathsf{ef}}$} (subsection 8.1). Then we comment on the nature of the universal system (subsection 8.2).
8.1. The universality result
As mentioned, with our preparation, the universality result will follow from the category-theoretic Fraïssé theorem after establishing the joint embedding property and the amalgamation property. We will first define this terminology, then establish these properties, and finally get the universality result.
We recall the definitions (from [32]), illustrated in figure 2. Let $\mathsf{C}$ be a category in which all morphisms are monic and let $\mathsf{D}$ be a full subcategory. An object $U$ in $\mathsf{C}$ is $\mathsf{D}$-universal, if, for any object $A$ in $\mathsf{D}$, there is a morphism $f:A\to U$ in $\mathsf{C}$.999[66, 1757] instead speaks of a ‘cofinal’ object, to avoid the category-theoretic connotation that a universal object is canonically determined by a universal property. An object $U$ in $\mathsf{C}$ is $\mathsf{D}$-homogeneous, if, for any object $A$ in $\mathsf{D}$ and morphisms $f,g:A\to U$ in $\mathsf{C}$, there is an isomorphism $h:U\to U$ in $\mathsf{C}$ such that $h\circ g=f$. We say $\mathsf{D}$ has the joint embedding property if for any objects $A$ and $B$ in $\mathsf{D}$, there is an object $C$ in $\mathsf{D}$ with morphisms $f:A\to C$ and $g:B\to C$ in $\mathsf{D}$. We say $\mathsf{D}$ has the amalgamation property if for any objects $A$, $B$, and $B^{\prime}$ in $\mathsf{D}$ and morphisms $f:A\to B$ and $f^{\prime}:A\to B^{\prime}$ in $\mathsf{D}$, there is an object $C$ in $\mathsf{D}$ with morphisms $g:B\to C$ and $g^{\prime}:B^{\prime}\to C$ in $\mathsf{D}$ such that $g\circ f=g^{\prime}\circ f^{\prime}$.
[…]
(a) Universal
[…]
(b) Homogeneous
[…]
(c) Joint embedding
[…]
(d) Amalgamation
Now the category-theoretic Fraïssé theorem of [32, thm. 1.1] says:
Theorem 8.1 ([32]).
We now apply this theorem to $\mathsf{C}={\mathsf{Sys}^{\mathsf{ef}}}$, which we have shown to be algebroidal. Thus, proving our universality result reduces to showing that the full subcategory $\mathsf{C}_{\mathsf{fin}}=\mathsf{Sys}^{\mathsf{ef}}_{\mathsf{fin}}$ of category-theoretically objects has the joint embedding property and the amalgamation property. Fortunately, here this is straightforward.
@@ 8.1. The universality result
Now, our main result follows using the Fraïssé-limit (theorem 8.1).
Corollary 8.4.
There is a system $(U,T)$ in $\mathsf{Sys}^{\mathsf{ef}}$} that is:
  1. (1)
    Universal: for any system $(Y,S)$ in $\mathsf{Sys}^{\mathsf{ef}}$}, there is an ef-pair $(e,f)$ from $(Y,S)$ to $(U,T)$.
  2. (2)
    Homogeneous: If $(Y,S)$ is a system in $\mathsf{Sys}^{\mathsf{ef}}$} with a finite state space and $(e,f)$ and $(e^{\prime},f^{\prime})$ are two ef-pairs from $(Y,S)$ to $(U,T)$, then there is an isomorphism $(\varphi,\psi):(U,T)\to(U,T)$ such that $(e,f)=(\varphi,\psi)\circ(e^{\prime},f^{\prime})$.
As already mentioned in section 2, from a dynamical systems perspective, we consider directed edges $x\xrightarrow{T}x^{\prime}$, rather than undirected (i.e., reflexive and symmetric) edges. But, since theorem 6.3 applies to any subcategory of $\mathsf{Sys}^{\mathsf{ef}}$}, we can, as mentioned in section 2, also recover the existence result of the projectively universal homogeneous (undirected) graph [57, 16, 40]. Specifically, call a system $(X,T)$ in $\mathsf{Sys}^{\mathsf{ef}}$} a graph if, for all $x,x^{\prime}\in X$, we have $x\xrightarrow{T}x$ (reflexivity) and, if $x\xrightarrow{T}x^{\prime}$, then $x^{\prime}\xrightarrow{T}x$ (symmetry). Let $\mathsf{G}^{\mathsf{ef}}$ be the full subcategory of $\mathsf{Sys}^{\mathsf{ef}}$} consisting of those systems that are graphs. Then conditions 1 and 2 of theorem 6.3 are readily checked, so $\mathsf{G}^{\mathsf{ef}}$ is algebroidal and the full subcategory $\mathsf{G}_{\mathsf{fin}}^{\mathsf{ef}}$ of category-theoretically finite objects consists of precisely the finite graphs. The proofs of propositions 8.2 and 8.3 still work and yield the amalgamation property and the joint embedding property for $\mathsf{G}_{\mathsf{fin}}^{\mathsf{ef}}$. Hence the Fraïssé limit theorem yields that $\mathsf{G}^{\mathsf{ef}}$ has a $\mathsf{G}^{\mathsf{ef}}$-universal and $\mathsf{G}^{\mathsf{ef}}_{\mathsf{fin}}$-homogeneous system.
@@ 8.2. The nature of the universal system
We end this section with some comments about the nature of this universal dynamical system $(U,T)$: its symmetry, universality, complexity, and its status as an analog computer.
Symmetry. The homogeneity property, that makes $(U,T)$ unique, is quite remarkable. It says that $(U,T)$ is ‘maximally symmetric’: every way of embedding a finite dynamical system into the universal system corresponds to a symmetry (i.e. isomorphism) of the universal system. Thus, the embedding of each finite system into the universal one is, in a sense, unique up to isomorphism.
@@
    Since every Turing machine can be seen as a dynamical system, it is a factor of and embeds into the universal system. Such an ef-pair can be seen as a simulation: each state of the Turing machine corresponds to a set of states of the universal system, and each dynamic step of the Turing machine is matched by a step in the universal system, and vice versa. Thus, in this sense, $(U,T)$ is Turing universal.
@@
Complexity. One would expect the universal system to be a very ‘complicated’ or ‘complex’ object since it contains every other system. However, at least its dynamics is in fact very simple: its orbits consist of at most two states. This surprising fact is known for the universal profinite graph, proven by [16, prop. 12] (also discussed by [40]). However, in our systems setting, a much simpler proof is available.
To state it, let’s fix some terminology. Let $(X,T)$ be a system. An orbit in $(X,T)$ is a sequence $(x_{n})_{n=0}^{\kappa}$ in $X$ with $\kappa\leq\omega$ and $x_{n}\xrightarrow{T}x_{n+1}$ for all $n<\kappa$. So the orbit is finite if $\kappa<\omega$ or infinite if $\kappa=\omega$. If $x\xrightarrow{T}y$, we call $y$ a successor state of state $x$. We call a state $x\in X$ a dead-end (or sink) if $x$ has no successor states. A loop in $(X,T)$ is a finite orbit $(x_{0},\ldots,x_{\kappa})$ with $x_{0}=x_{\kappa}$ and $\kappa\geq 1$. It is a self-loop if $\kappa=1$.
@@
Orbits in the universal system $(U,T)$ have at most $2$ states. In other words, for every state $x\in U$, either $x$ is a dead-end or all its successor states are dead-ends. In particular, $(U,T)$ does not have any loops (also no self-loops).
@@
Toward a contradiction, assume there was an orbit with three states: $x_{0}\xrightarrow{T}x_{1}\xrightarrow{T}x_{2}$. Consider the system $(Y,S)$ with $Y=\{y_{0},y_{1}\}$ and $S(y_{0})=\{y_{1}\}$ and $S(y_{1})=\emptyset$. It is a non-trivial system and hence is in $\mathsf{Sys}^{\mathsf{ef}}$}. So there is a factor $f:(U,T)\to(Y,S)$. We cannot have $f(x_{0})=y_{1}$: Otherwise, since $f$ is a factor and $x_{0}\xrightarrow{T}x_{1}$, there must be $y^{\prime}\in Y$ with $f(x_{1})=y^{\prime}$ and $y_{1}\xrightarrow{S}y^{\prime}$, but $y_{1}$ is a dead-end. Hence $f(x_{0})=y_{0}$. Similarly, we cannot have $f(x_{1})=y_{1}$. Hence $f(x_{1})=y_{0}$. But then, since $x_{0}\xrightarrow{T}x_{1}$ and $f$ is a factor, we have $f(x_{0})\xrightarrow{S}f(x_{1})$, and hence $y_{0}\xrightarrow{S}y_{0}$, contradiction. ∎
Analog computer? In section 2, we said that every analog computer can be seen as a dynamical system. But we also noted that the converse is a difficult philosophical question. Accordingly, it is out of scope of this paper to discuss whether the universal system $(U,T)$ is an analog computer, and hence is a universal analog computer. But we pick this up again as an interesting open question in section 12.
@@ 9.2. No universal deterministic system
    [23, Corollary 4.2]: If $(X,T)$ is a compact metric system, then there is a Cantor system $(2^{\omega},S)$ and a factor $f:(2^{\omega},S)\to(X,T)$. (This extends the universality of Cantor space from spaces to systems.)
@@
    [23, Remark 4.6]: Assume $Z$ is a compact system that is universal for compact metric systems, i.e., if $(X,T)$ is a compact metric system, there is a factor $f:(Z,R)\to(X,T)$. Then $Z$ cannot be metrizable. (In fact, $(Z,R)$ cannot even factor onto all rotations of the circle.)
Toward a contradiction, assume $(Z,R)$ is universal in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$. Then $(Z,R)$ is a compact system, and it is universal for compact metric systems: If $(X,T)$ is a compact metric system, then, by 1, there is a Cantor system $(2^{\omega},S)$ and a factor $f:(2^{\omega},S)\to(X,T)$. Hence $(2^{\omega},S)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$, so, by universality, there is a factor $g:(Z,R)\to(2^{\omega},S)$. In sum, there is a factor $f\circ g:(Z,R)\to(X,T)$, as needed. However, then 2 implies that $Z$ is not metrizable, in contradiction to $Z$ being a Polish space. ∎
@@
Toward a contradiction, assume there is a left adjoint (aka reflector) $\mathsf{R}:{\mathsf{Sys}^{\mathsf{ef}}}\to\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ with unit $\eta$. Let $(U,T)$ be the universal system in $\mathsf{Sys}^{\mathsf{ef}}$} (from corollary 8.4). We claim that $\mathsf{R}(U,T)$ is universal in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$, hence contradicting proposition 9.2. Indeed, let $(Y,S)$ be in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$. We need to find an ef-pair $(Y,S)\to\mathsf{R}(U,T)$. By universality of $(U,T)$, there is an ef-pair $(e,f):(Y,S)\to(U,T)$. Hence, qua left adjoint, there is a unique ef-pair $\mathsf{R}(e,f)=(g,h)$ making the following diagram commute:
@@ 9.3. The category of deterministic systems still is semi-algebroidal
We start by recalling some terminology from dynamical systems theory [71]. Let $A$ be a finite set, called alphabet. The (one-sided) full shift over $A$ is the system $(A^{\omega},\sigma)$ where $A^{\omega}$ carries the product topology of the discrete topology on $A$ and $\sigma$ is the shift map, i.e., $\sigma(x)_{n}=x_{n+1}$. A shift (aka shift space) is a closed, shift-invariant subset $X$ of a full shift on some alphabet. Hence, if $X$ is nonempty, $(X,\sigma)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$.
A topological generator of a system $(X,T)$ in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is a closed partition $\mathcal{C}$ of $X$ (i.e., a partition of $X$ whose partition cells are closed subsets of $X$) such that $\{T^{-n}(C):C\in\mathcal{C},n\in\omega\}$ generates the topology of XX. [^12] Recall that a collection $\mathcal{F}$ of subsets of a topological space $X$ generates the topology $\tau$ of $X$ if $\tau$ is the smallest topology containing all subsets of $\mathcal{F}$. We say $(X,T)$ has a finite topological generator if there is a finite clopen partition of $X$ that is a topological generator of $X$.
If $(X,T)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ and $\mathcal{C}$ is a finite clopen partition of $X$, the $\mathcal{C}$-name of $x\in X$ is the sequence $(C_{n})$ of $\mathcal{C}$-partition cells where, for each $n$, $C_{n}$ is the cell with $T^{n}(x)\in C_{n}$. Then the set $Y$ of all $\mathcal{C}$-names is a shift over alphabet $\mathcal{C}$ and the function $f:X\to Y$ that maps $x$ to its $\mathcal{C}$-name is a factor. We call $Y$ the $\mathcal{C}$-shift and $f$ the $\mathcal{C}$-factor.
@@
($\Leftarrow$) Now assume that $\mathsf{DS}^{\mathsf{ef}}$ satisfies requirement 2 and let $(X,T)$ be category-theoretically finite. So $(X,T)$ is a colimit of an $\omega$-chain $((X_{i},T_{i}),(e_{ii+1},f_{ii+1}))$ of shifts in $\mathsf{DS}^{\mathsf{ef}}$. By lemma 6.1, $(X,T)$ is isomorphic to some $(X_{i},T_{i})$, which, qua shift, has a topological generator (proposition 9.5). ∎

### v2

@@ 6.2. Proof of the algebroidality theorem
The remaining part (2) of being semi-algebroidal we show in two steps. First, we show it when understanding ‘finite’ in the set-theoretic sense, and then we show that this is equivalent to the category-theoretic sense.
We start, though, with a lemma, which is a version of [64, lem. 2.1]. For completeness, we add a proof in appendix A.3.
@@
To show the equivalence of the two senses of ‘finite’, we need two more lemmas. The first is a useful fact about algebroidal categories.
Lemma 6.9.
Let $\mathsf{C}$ be a category in which all morphisms are monic. Let $B$ be a category-theoretically finite object of $\mathsf{C}$. If $(B,f_{i})$ is the colimit of an $\omega$-chain $(A_{i},f_{ii+1})_{i\in\omega}$ in $\mathsf{C}$, then some $f_{i}:A_{i}\to B$ is an isomorphism.
Proof.
The second lemma is a standard fact about the interplay between partitions and projections in inverse limits. For completeness, we add a proof in appendix A.3.
@@ 7. The universal and homogeneous nondeterministic system
Now we can show the universality result for nondeterministic systems (subsection 7.1). Then we comment on the nature of the universal system (subsection 7.2).
7.1. Proof of theorem 4.2
Having established the algebroidality of $\mathsf{Sys}^{\mathsf{ef}}$}, it only remains, according to our proof strategy (section 4.2), to establish the joint embedding property and the amalgamation property of $\mathsf{Sys}^{\mathsf{ef}}_{\mathsf{fin}}$. (To recall, $\mathsf{Sys}^{\mathsf{ef}}_{\mathsf{fin}}$ is the full subcategory of $\mathsf{Sys}^{\mathsf{ef}}$} consisting of the category-theoretically finite objects in $\mathsf{Sys}^{\mathsf{ef}}$}.) Then theorem 4.2 follows from the Fraïssé theorem. Fortunately, here this is straightforward.
@@
This finishes the proof of theorem 4.2. As noted, the theorem generalizes the result for profinite undirected graphs. More precisely:
Remark 7.3.
As mentioned in section 2, from a dynamical systems perspective, we consider directed edges $x\xrightarrow{T}x^{\prime}$, rather than undirected (i.e., reflexive and symmetric) edges. But, since theorem 6.2 applies to any subcategory of $\mathsf{Sys}^{\mathsf{ef}}$}, we can, as mentioned in section 2, also recover the existence result of the projectively universal homogeneous (undirected) graph [64, 20, 47]. Specifically, call a system $(X,T)$ in $\mathsf{Sys}^{\mathsf{ef}}$} a graph if, for all $x,x^{\prime}\in X$, we have $x\xrightarrow{T}x$ (reflexivity) and, if $x\xrightarrow{T}x^{\prime}$, then $x^{\prime}\xrightarrow{T}x$ (symmetry). Let $\mathsf{G}^{\mathsf{ef}}$ be the full subcategory of $\mathsf{Sys}^{\mathsf{ef}}$} consisting of those systems that are graphs. Then conditions 1 and 2 of theorem 6.2 are readily checked, so $\mathsf{G}^{\mathsf{ef}}$ is algebroidal and the full subcategory $\mathsf{G}_{\mathsf{fin}}^{\mathsf{ef}}$ of category-theoretically finite objects consists of precisely the finite graphs. The proofs of propositions 7.1 and 7.2 still work and yield the amalgamation property and the joint embedding property for $\mathsf{G}_{\mathsf{fin}}^{\mathsf{ef}}$. Hence the Fraïssé limit theorem yields that $\mathsf{G}^{\mathsf{ef}}$ has a $\mathsf{G}^{\mathsf{ef}}$-universal and $\mathsf{G}^{\mathsf{ef}}_{\mathsf{fin}}$-homogeneous system.
@@ 7.2. The nature of the universal system
We end this section with some comments about the nature of the universal nondeterministic system $(U,T)$: its symmetry, universality, complexity, and its status as an analog computer.
Symmetry. The homogeneity property, that makes $(U,T)$ unique, is quite remarkable. In section 4.1, we motivated it as saying that ‘every part of $(U,T)$ looks the same’. Another way of saying this is that $(U,T)$ is ‘maximally symmetric’: whenever we have two ways of embedding a finite dynamical system into the universal system, there is a symmetry (i.e. isomorphism) of the universal system that transforms one embedding into the other. Thus, the embedding of each finite system into the universal one is, in a sense, unique up to isomorphism.
@@
    Since every Turing machine can be seen as a dynamical system, it is a factor of and embeds into the universal system. Such an ef-pair can be seen as a simulation (see remark 3.10). Thus, in this sense, $(U,T)$ is Turing universal.
@@
Complexity. One would expect the universal system to be a very ‘complicated’ or ‘complex’ object since it contains every other system. However, it actually is surprisingly simple, as we will show now: Its state space is a Cantor space (i.e., isomorphic to $2^{\omega}$), and its dynamics has no complicated long-term behavior because the orbits consist of at most two states. The latter, surprising fact is known for the universal profinite graph, proven by [20, prop. 12] (also discussed by [47]). However, in our systems setting, a much simpler proof is available.
First, the claim about the state space [[, cf.]13]Geschke2022.
@@
The state space of the universal system $(U,T)$ from theorem 4.2 is a Cantor space (i.e., isomorphic to $2^{\omega}$).
@@
By Brouwer’s theorem [[, e.g.]thm. 7.4]Kechris1995, a space is a Cantor space if it is a nonempty, compact, zero-dimensional, metrizable space without isolated points (i.e., no points whose singleton is open). Since $(U,T)$ is in $\mathsf{Sys}^{\mathsf{ef}}$}, it hence remains to show that $U$ has no isolated points.
Toward a contradiction, assume $x_{0}\in U$ is isolated. Consider the two-element discrete space $Y=\{y_{0},y_{1}\}$ and the function $f:U\to Y$ which maps $x_{0}$ to $y_{0}$ and any $x\in U\setminus\{x_{0}\}$ to $y_{1}$—so $f$ is a continuous function. Define the dynamics $S$ on $Y$ by: $y\xrightarrow{S}y^{\prime}$ iff there is $x,x^{\prime}\in U$ with $f(x)=y$, $f(x^{\prime})=y^{\prime}$, and $x\xrightarrow{T}x^{\prime}$. So $f:(U,T)\to(Y,S)$ is a factor. Now define the system $(Z,R)$ by ‘doubling’ $y_{0}$: let $Z:=\{y_{0},z_{0},y_{1}\}$ be a three-element discrete space and let $R(z):=S(z)$ if $z\in\{y_{0},y_{1}\}$ and $R(z_{0}):=\emptyset$. Define $g:Z\to Y$ by mapping $y_{0}\mapsto y_{0}$, $z_{0}\mapsto y_{0}$, and $y_{1}\mapsto y_{1}$. So $g:(Z,R)\to(Y,S)$ is a factor. By universality, there is a factor $h:(U,T)\to(Z,R)$, and, by homogeneity, we can choose $h$ such that $g\circ h=f$. [^17] This is the idea of $(U,T)$ being a ‘saturated object’, which we will define and use again in the proof of proposition 10.4. Hence, since $g(h(x_{0}))=f(x_{0})=y_{0}$ and $g^{-1}(y_{0})=\{y_{0},z_{0}\}$, we must have either $h(x_{0})=y_{0}$ or $h(x_{0})=z_{0}$. Let’s consider the first case (the other is analogous). By surjectivity, there is $x\in U$ with $h(x)=z_{0}$. But then $x\in U\setminus\{x_{0}\}$, so $y_{1}=f(x)=g(h(x))=g(z_{0})=y_{0}$, contradiction. ∎
Now, the claim about the dynamics. Let us first fix some terminology. Let $(X,T)$ be a system. An orbit in $(X,T)$ is a sequence $(x_{n})_{n=0}^{\kappa}$ in $X$ with $\kappa\leq\omega$ and $x_{n}\xrightarrow{T}x_{n+1}$ for all $n<\kappa$. So the orbit is finite if $\kappa<\omega$ or infinite if $\kappa=\omega$. If $x\xrightarrow{T}y$, we call $y$ a successor state of state $x$. We call a state $x\in X$ a dead end (or sink) if $x$ has no successor states. A loop in $(X,T)$ is a finite orbit $(x_{0},\ldots,x_{\kappa})$ with $x_{0}=x_{\kappa}$ and $\kappa\geq 1$. It is a self-loop if $\kappa=1$.
Proposition 7.5.
Orbits in the universal system $(U,T)$ have at most $2$ states. In other words, for every state $x\in U$, either $x$ is a dead end or all its successor states are dead ends. In particular, $(U,T)$ does not have any loops (also no self-loops).
Proof.
Toward a contradiction, assume there was an orbit with three states: $x_{0}\xrightarrow{T}x_{1}\xrightarrow{T}x_{2}$. Consider the system $(Y,S)$ with $Y=\{y_{0},y_{1}\}$ and $S(y_{0})=\{y_{1}\}$ and $S(y_{1})=\emptyset$. It is a non-trivial system and hence is in $\mathsf{Sys}^{\mathsf{ef}}$}. So there is a factor $f:(U,T)\to(Y,S)$. We cannot have $f(x_{0})=y_{1}$: Otherwise, since $f$ is a factor and $x_{0}\xrightarrow{T}x_{1}$, there must be $y^{\prime}\in Y$ with $f(x_{1})=y^{\prime}$ and $y_{1}\xrightarrow{S}y^{\prime}$, but $y_{1}$ is a dead end. Hence $f(x_{0})=y_{0}$. Similarly, we cannot have $f(x_{1})=y_{1}$. Hence $f(x_{1})=y_{0}$. But then, since $x_{0}\xrightarrow{T}x_{1}$ and $f$ is a factor, we have $f(x_{0})\xrightarrow{S}f(x_{1})$, and hence $y_{0}\xrightarrow{S}y_{0}$, contradiction. ∎
Analog computer? In section 2, we said that every analog computer can be seen as a dynamical system. But we also noted that the converse is a difficult philosophical question. Accordingly, it is out of scope of this paper to discuss whether the universal system $(U,T)$ is an analog computer, and hence is a universal analog computer. We only note that the simple nature of $(U,T)$ seems to intuitively speak in favor of $(U,T)$ being an analog computer—if it were an extremely intricate and complicated system, it would seem less plausible to view it as a ‘computing machine’. We pick this up again as an interesting open question in section 11.
@@ 8.2. No universal deterministic system
    [29, Corollary 4.2]: If $(X,T)$ is a compact metric system, then there is a Cantor system $(2^{\omega},S)$ and a factor $f:(2^{\omega},S)\to(X,T)$. (This extends the universality of Cantor space from spaces to systems.)
@@
    [29, Remark 4.6]: Assume $Z$ is a compact system that is universal for compact metric systems, i.e., if $(X,T)$ is a compact metric system, there is a factor $f:(Z,R)\to(X,T)$. Then $Z$ cannot be metrizable. (In fact, $(Z,R)$ cannot even factor onto all rotations of the circle.)
Toward a contradiction, assume $(Z,R)$ is universal in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$. Then $(Z,R)$ is a compact system, and it is universal for compact metric systems (not just zero-dimensional ones): If $(X,T)$ is a compact metric system, then, by 1, there is a Cantor system $(2^{\omega},S)$ and a factor $f:(2^{\omega},S)\to(X,T)$. Hence $(2^{\omega},S)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$, so, by universality, there is a factor $g:(Z,R)\to(2^{\omega},S)$. In sum, there is a factor $f\circ g:(Z,R)\to(X,T)$, as needed. However, then 2 implies that $Z$ is not metrizable, in contradiction to $Z$ being a Polish space. ∎
@@
Toward a contradiction, assume there is a left adjoint (aka reflector) $\mathsf{R}:{\mathsf{Sys}^{\mathsf{ef}}}\to\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ with unit $\eta$. Let $(U,T)$ be the universal system in $\mathsf{Sys}^{\mathsf{ef}}$} (from theorem 4.2). We claim that $\mathsf{R}(U,T)$ is universal in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$, hence contradicting proposition 8.2. Indeed, let $(Y,S)$ be in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$. We need to find an ef-pair $(Y,S)\to\mathsf{R}(U,T)$. By universality of $(U,T)$, there is an ef-pair $(e,f):(Y,S)\to(U,T)$. Hence, qua left adjoint, there is a unique ef-pair $\mathsf{R}(e,f)=(g,h)$ making the following diagram commute:
@@ 8.3. The category of deterministic systems still is semi-algebroidal
We start by recalling some terminology from dynamical systems theory [81]. A finite topological generator of a system $(X,T)$ in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is a finite clopen partition $\mathcal{C}$ of $X$ (i.e., $\mathcal{C}$ has finitely many cells and they are clopen) such that $\{T^{-n}(C):C\in\mathcal{C},n\in\omega\}$ generates the topology of XX. [^19] Recall that a collection $\mathcal{F}$ of subsets of a topological space $X$ generates the topology $\tau$ of $X$ if $\tau$ is the smallest topology containing all subsets of $\mathcal{F}$. If $(X,T)$ is in $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ and $\mathcal{C}$ is a finite clopen partition of $X$, the $\mathcal{C}$-name of $x\in X$ is the sequence $(C_{n})$ of $\mathcal{C}$-partition cells where, for each $n$, $C_{n}$ is the cell with $T^{n}(x)\in C_{n}$. Then the set $Y$ of all $\mathcal{C}$-names is a shift over alphabet $\mathcal{C}$ and the function $f:X\to Y$ that maps $x$ to its $\mathcal{C}$-name is a factor. We call $Y$ the $\mathcal{C}$-shift and $f$ the $\mathcal{C}$-factor.
@@
($\Leftarrow$) Now assume that $\mathsf{DS}^{\mathsf{ef}}$ satisfies requirement 2 and let $(X,T)$ be category-theoretically finite. So $(X,T)$ is a colimit of an $\omega$-chain $((X_{i},T_{i}),(e_{ii+1},f_{ii+1}))$ of shifts in $\mathsf{DS}^{\mathsf{ef}}$. By lemma 6.9, $(X,T)$ is isomorphic to some $(X_{i},T_{i})$, which, qua shift, has a finite topological generator (proposition 8.5). ∎